# Women are less comfortable expressing opinions online than men and report heightened fears for safety: Surveying gender differences in experiences of online harms

### v1

@@ 2.3 Survey
Demographics and background questions: For each participant, we collected standard demographic information about age, gender, ethnicity, education and political orientation. Age could be entered as any number with a minimum of 18. For gender, ethnicity and education level, participants were asked to select the option that they felt best described them from a list of standard predefined categories. For political orientation, participants were asked to select the option that they felt most described their beliefs: ‘More to the left’, ‘Centre’, ‘More to the right’ and ‘Prefer not to say’. Participants were also asked what kind of device they were using to complete the survey, as well as how often they use the internet, whether they use/have used social media before, and how often. All demographic questions, other than age, provided participants with a ‘prefer not to say’ option (this was not included for age as being 18 or over was a requirement to participate).
Exposure to online harms: We presented participants with a list of 15 online harms along with definitions for each. The harms we included were: Hate speech, Misinformation, Misogyny, Trolling, Bullying, Cyberstalking, Cyberflashing, Group attacks, Impersonation, Catfishing, Threats of non-sexual violence, Threats of sexual violence, Doxing, Image based sexual abuse, and Eating disorder content. Participants were asked to indicate whether they had heard of or seen each harm online in the past year (Never heard of/ Heard of but not seen in the last year/ Seen online in the last year).
For each harm, if participants indicated they had seen it in the past year, they were then asked a follow up about the extent to which they had directly received each harm and the extent to which they had witnessed each harm. For ‘content-based’ harms (hate speech, misinformation, misogyny and eating disorder content), ‘directly received’ meant the content was directly intended for the participant, for example naming them or being sent in a direct message, while for ‘contact-based’ harms (trolling, bullying, stalking, cyberflashing, group attacks, catfishing, threats of physical violence, threats of sexual violence, doxing and image-based sexual abuse), ‘directly received’ meant the harm happened to them. In the case of impersonation, directly received meant that someone impersonated the participant’s account. ‘Witnessing’ always meant that they had seen the content but that it was not directly intended for them personally. Response options for both scales were: Never/ Once/ 2-4 times/ 5 or more times/ Prefer not to say.
@@
Fears of exposure to online harms: To measure fears about exposure to each of the same 15 harms, participants were asked to indicate the extent to which they fear witnessing such content, and the extent to which they fear directly receiving such content (both scales: Not at all / Not very much / Somewhat / Very much).
@@
Comfort with online participation: Participants were asked how comfortable they are (Not at all/ Not very/ Somewhat/ Extremely) engaging with seven different online activities in both public/open settings and private/restricted settings. The behaviours we asked about were: Expressing political opinions online; Expressing other opinions online (e.g., opinions about the news, TV shows or music); Challenging content they see online and do not agree with; Sharing personal information online (such as name, date of birth, where they live and work, gender, religious beliefs, and so on); Sharing photos of themselves online; Sharing photos of friends and family online; Sharing photos of activities online (e.g., travelling, cooking or going out with friends).
Psychological impact: Participants were asked to indicate the extent to which certain experiences online had caused them particular kinds of feelings. These six feelings were: Feeling sad or low; Feeling angry or frustrated; With physical symptoms (such as insomnia, headaches and stomach aches); Feeling as though their job/career had been negatively affected; Feeling as though relationships with a partner, friends or family had been negatively affected; Feeling (more) eager or keen to use social media to advocate for a specific cause, political ideology or to educate others on a specific topic (Response scales for all were: Not at all/ Not very much/ Somewhat/ Very much).
Engagement with safety tools: Participants were asked to indicate if they had ever used seven types of online safety features on social media platforms. These were: Disabled location sharing on a device; Disabled airdrop and/or Bluetooth on a device; Made profile/account/page private; Limited who can contact you (for example so you cannot receive messages from people you don’t know); Limited who can engage with your contents (for example who can like or comment on your posts); Limited who can tag/mention you in images/posts/tweets; Limited how people can find your profile (for example controlling whether people can find your profile by linking with email or phone, through a search engine or by appearing in suggested people lists) (Yes/No for all).
@@ 2.4 Procedure
After participants gave their informed consent to take part in the survey, they responded to the demographic questions, questions about device use, time spent online, and social media use. Following this, they read brief definitions of the fifteen harms so as to familiarise themselves with what each one means. They were asked to pay close attention to the definitions and could move on after a minimum of 45 seconds. Participants then responded to the questions about harm exposure for each of the fifteen harms (awareness/exposure, followed by extent of exposure and perpetrator identity if relevant), followed by questions about fears (including fear origins and subsequent behavioural changes if relevant). Next, participants answered questions about comfort with various forms of online participation, followed by questions about psychological impact of online experiences, followed by questions about engagement with safety tools online. At the end of the questions, participants were given an opportunity to provide feedback in a free text box before continuing to the debrief and finally completing the submission and being returned to Prolific for payment. The survey was designed to take approximately 15 minutes to complete and each participant received £2.25 for their time.
@@ 2.5 Data analysis
Additionally, we test for associations between fears about online harms and comfort engaging in several forms of online participation by running a series of 84 logistic regression analyses, each including fear as predictor (one for fear about each of 12 harms, see Results for harms included), and comfort engaging in the online behaviour of interest as the outcome (one for comfort with each of the seven behaviours), all controlling for gender, age, and internet use. For all statistical analyses, the accepted significance threshold was set at .05.
@@ 3.1.1 Awareness and exposure
Gender differences in exposure to various harms show mixed results. Women are 22% more likely to say they have seen online misogyny (p=.027), 24% more likely to say they have seen bullying (p=.017) and 104% more likely to say they have seen eating disorder content (p<<<.001) than men. However, women are 20% less likely to say they have seen hate speech (p=.016), 25% less likely to say they have seen misinformation (p=.017), 23% less likely to say they have seen seen impersonation (p=.007), 31% less likely to say they have seen threats of physical violence (p<<<.001), and 29% less likely to say they have seen doxing (p=.003) than men. For many of the harms, such as trolling and cyberstalking, differences in self-reported exposure between men and women are not statistically significant.
@@
Image-based sexual abuse Female 8.4%
@@
Eating disorder content Female 27.1% ***
@@
Table 1: Proportions of men and women indicating they have seen each harm in the past year. In the Sig. column, * means significant at the p<<<.05 level; ** at the p<<<.01 level, and *** at the p<<<.001 level.
@@ 3.1.2 Extent of exposure
Importantly, the results on exposure presented above show whether or not people have seen certain types of content but not whether they have been direct targets of online harms, nor how much. To understand the extent of people’s exposure more deeply, for each of the 15 harms that participants indicated they had seen in the past year, we analysed responses to follow up questions about extent to which they had witnessed such content and the extent to which they had directly received such content (see Methods). Response options for both scales were converted into 4 point numeric scales (1 = Never; 2 = Once; 3 = 2-4 times; 4 = 5 or more times, with ‘Prefer not to say’ as NA) and non-parametric t-tests (Mann-Whitney U-Tests) examined gender differences in being directly targeted by each type of harm. [^3] An alternative analysis is to create binary responses (Never = 0 / At least once = 1) and use logistic regressions to examine the effect of gender on whether people have been targeted before or not. Taking this approach yields identical results. If participants had previously indicated that they had not seen the harm in the past year and were therefore not asked follow-up questions about the capacity in which they had seen it, their responses for this question were re-coded from missing to ’Never’.
@@
Women reported being the direct targets of several online harms to a significantly greater extent than men, including online misogyny (p<<<.001), cyberstalking (p=.026), cyberflashing (p<<<.001), eating disorder content (p=.039) and, marginally, image based abuse (p=.075). Table 4 in Supplementary Information shows proportions of men and women choosing each response option for the five harms that women report being targeted by to a greater extent than men. Men reported being the direct targets of hate speech (p<<<.001), misinformation (p<<<.001), trolling (p<<<.001) and threats of physical violence (p=.007) to a greater extent than women. Table 5 in Supplementary Information shows proportions of men and women choosing each response option for the four harms that men report being targeted by to a greater extent than women. There were no significant gender differences in reports of being directly targeted by the remaining six harms. Figure 1 shows the extent to which men and women report having been direct targets of each online harm. Note that while we asked individuals who had been targeted by harms whether they knew the identity(ies) of the perpetrator(s), comparing proportions of responses led to sample sizes too small to make meaningful comparisons for this question. However, we share these responses in our open dataset.
@@
Figure 1: The extent to which people report having been direct targets of each online harm by gender. F>>>M indicates women are targeted significantly more than men. M<<<F indicates men are targeted significantly more than women. ‘Prefer not to say’ responses were uncommon and are not included.
@@
Comparing gender differences in witnessing each type of harm (an individual seeing harmful content online not directly targeted at them), men reported witnessing hate speech, misinformation, trolling, impersonation, catfishing, threats of physical violence, threats of sexual violence, doxing, and image-based abuse to a greater extent than women (all ps<<<.05), while women reported witnessing eating disorder content (p<<<.001) and, marginally, misogyny (p=.051) and bullying (p=.073) to a greater extent than men. [^4] We explored whether possible gender differences in overall internet use explained these results but we found men and women report using the internet and social media to a similar extent overall. The differences here in results for witnessing and directly receiving online harms shows the importance of understanding both who is most likely to be targeted, alongside understanding overall exposure.
@@ 3.2 Psychological impact of online experiences
We asked participants the extent to which experiences online had ever caused them six particular kinds of feelings (see Methods). For each of these responses, we created a binary outcome variable to indicate whether they had experienced this psychological effect or not (Yes, 1 - Somewhat/Very much and No, 0 - Not at all/Not very much). Overall, 48% of respondents said they had been left feeling sad or low as a result of an online experience; 68% that they had felt angry or frustrated; 11% with physical symptoms (such as insomnia, headaches and stomach aches); 9% as though their career had been negatively affected; 15% as though their relationships had been negatively affected; and 22% feeling (more) eager or keen to use social media to advocate for a specific cause. Logistic regression analyses with gender as predictor and response as outcome for each of the items found that women were significantly more likely than men to have been left feeling sad or low (96% more likely), angry or frustrated (54% more likely), and with physical symptoms (47% more likely) as a result of an experience online (all ps<<<.001). There were no gender differences in responses for the remaining three questions about psychological impact (all ps >>>.05). Table 6 in Supplementary Information shows responses to each item for men and women.
@@ 3.3.1 Overall extent of fears about receiving online harms
To compare gender differences in fears about exposure to each of the 15 online harms, we created a binary outcome for fear, with ‘Somewhat’ and ‘Very much’ as 1 - Fearful, and ‘Not at all’ and ‘Not very much’ as 0 - Not fearful. Women consistently express significantly greater levels of fear than men across all 15 types of content, both for witnessing and for directly receiving each harm. Women are 117% more likely to fear receiving hate speech than men, 39% more likely to fear receiving misinformation, 489% more likely to fear receiving misogyny, 125% more likely to fear being targeted by trolling and bullying, 120% more likely to fear being targeted by cyberstalking, 196% more likely to fear being targeted by cyberflashing, 68% more likely to fear being targeted by group attacks, 50% more likely to fear having an account impersonated, 64% more likely to fear catfishing, 112% more likely to fear receiving physically violent threats, 215% more likely to fear receiving sexually violent threats, 38% more likely to fear being targets of doxing, 133% more likely to fear being targets of image-based abuse, and 174% more likely to fear receiving eating disorder content (all ps<<<.001). Table 2 shows proportions of men and women fearing receiving each type of online harm.
@@
Image based sexual abuse Low fear 71.0% 51.2%
@@
Eating disorder content Low fear 87.3% 71.4%
@@ 3.3.2 Origins of fears about receiving online harms
When asked where they thought fears about receiving online harms might have come from, the top three choices for women were: The media (65%), Female friend’s experience (37%) and Public figure (35%). Additionally, 30% of women chose Personal experience, 8% chose Male friend’s experience, 5% Non-binary friend’s experience. The top three choices for men were: The media (60%), Personal experience (29%), and Female friend’s experience (27%). Additionally, 26% of men chose Public figure, 14% chose Male friend’s experience, and 3% Non-binary friend’s experience. Therefore, both men and women chose the media and female friends’ experiences as the most common sources of fears of online harms. [^5] We do not analyse free text responses here but these are available in our open dataset.
3.3.3 Behavioural effects of fears about receiving online harms

### v2

@@ 3.3 Survey
Demographics and background questions: For each participant, we collected standard demographic information about age, gender, ethnicity, education and political orientation. Age could be entered as any number with a minimum of 18. For gender, ethnicity and education level, participants were asked to select the option that they felt best described them from a list of standard predefined categories. For political orientation, participants were asked to select the option that they felt most described their beliefs: ‘More to the left’, ‘Centre’, ‘More to the right’ and ‘Prefer not to say’. Participants were also asked what kind of device they were using to complete the survey, as well as how often they use the internet, whether they use/have used social media before, and how often. All demographic questions, other than age, provided participants with a ‘prefer not to say’ option.
Exposure to online harms: We presented participants with a list of 15 online harms along with definitions for each. These harms were selected to be as comprehensive as possible and to reflect categories of interest in UK policy, drawing on lists developed by Ofcom and the Victims Commissioner (Ofcom,, 2022; Storry and Poppleton,, 2022). The harms we included were: Bullying, Catfishing, Cyberflashing, Cyberstalking, Doxing, Eating disorder promotion content, Group attacks, Hate speech, Image-based sexual abuse, Impersonation, Misogyny, Misinformation, Threats of non-sexual violence, Threats of sexual violence, and Trolling [^4] We note here that while the harms we included are defined separately and analysed as distinct categories, in practice they may overlap (for example, misinformation can contain hate, bullying can contain misogyny and so on). It is possible, then, that participants could have selected multiple harms to describe a single experience. Our analyses do not require the harms to be conceptually independent and the overlap across categories reflects the interconnected nature of online harms. (see Table 6.1.1 in Supplementary Information for the definitions we provided to participants). Following prior work (Enock et al.,, 2023), participants were asked to indicate whether they had heard of or seen each harm online in the past year (Never heard of/ Heard of but not seen in the last year/ Seen online in the last year).
For each harm, if participants indicated they had seen it in the past year, they were then asked a follow up about the extent to which they had directly received each harm and the extent to which they had witnessed each harm. For ‘content-based’ harms (hate speech, misinformation, misogyny and eating disorder promotion content), ‘directly received’ meant the content was directly intended for the participant, for example naming them or being sent in a direct message, while for ‘contact-based’ harms (trolling, bullying, stalking, cyberflashing, group attacks, catfishing, threats of physical violence, threats of sexual violence, doxing and image-based sexual abuse), ‘directly received’ meant the harm happened to them. In the case of impersonation, directly received meant that someone impersonated the participant’s account. ‘Witnessing’ always meant that they had seen the content but that it was not directly intended for them personally. Response options were: Never/ Once/ 2-4 times/ 5 or more times/ Prefer not to say.
@@
Fears of exposure to online harms: To measure fears about exposure to each of the same 15 harms, participants were asked to indicate the extent to which they fear witnessing such content, and the extent to which they fear directly receiving such content (both scales: Not at all / Not very much / Somewhat / Very much). This measure was adapted from standard approaches in the fear-of-crime literature, which often assess intensity of fear using similar Likert scales (e.g., Callanan and Teasdale,, 2009). We note here that we did not define fear for participants and so, consistent with the fear-of-crime literature, responses could reflect emotional reactions to the possibility of harm exposure, more cognitive risk assessments of such exposure, or a combination of the two.
@@
Comfort with online participation: Participants were asked how comfortable they are (Not at all/ Not very/ Somewhat/ Extremely) engaging with seven different online activities in both public/open settings and private/restricted settings. The behaviours we asked about were: Expressing political opinions online; Expressing other opinions online (e.g., opinions about the news, TV shows or music); Challenging content they see online and do not agree with; Sharing personal information online (such as name, date of birth, where they live and work, gender, religious beliefs, and so on); Sharing photos of themselves online; Sharing photos of friends and family online; Sharing photos of activities online (e.g., travelling, cooking or going out with friends). Items for the comfort measures were developed specifically for this study but were informed by prior work on online voice and participation Ofcom, (2022); Olson Carter and LaPoe, (2018); Papacharissi, (2015).
Psychological impact: Participants were asked to indicate the extent to which certain experiences online had caused them particular kinds of feelings. These six feelings were: Feeling sad or low; Feeling angry or frustrated; With physical symptoms (such as insomnia, headaches and stomach aches); Feeling as though their job/career had been negatively affected; Feeling as though relationships with a partner, friends or family had been negatively affected; Feeling (more) eager or keen to use social media to advocate for a specific cause, political ideology or to educate others on a specific topic (Response scales for all were: Not at all/ Not very much/ Somewhat/ Very much). We drew on prior research to inform the items we included (Amnesty International,, 2017; Stevens et al.,, 2021), but the set of six items used was developed specifically for this study.
Engagement with safety tools: Drawing on recent survey work examining uptake of online safety technology (Bright et al.,, 2024), participants were asked to indicate if they had ever used seven types of online safety features on social media platforms. These were: Disabled location sharing on a device; Disabled airdrop and/or Bluetooth on a device; Made profile/account/page private; Limited who can contact you (for example so you cannot receive messages from people you don’t know); Limited who can engage with your contents (for example who can like or comment on your posts); Limited who can tag/mention you in images/posts/tweets; Limited how people can find your profile (for example controlling whether people can find your profile by linking with email or phone, through a search engine or by appearing in suggested people lists) (Yes/No for all).
@@ 3.4 Procedure
After participants gave their informed consent to take part in the survey, they responded to the demographic questions, questions about device use, time spent online, and social media use. Following this, they read brief definitions of the fifteen harms so as to familiarise themselves with what each one means. They were asked to pay close attention to the definitions and could move on after a minimum of 45 seconds. Participants then responded to the questions about harm exposure for each of the fifteen harms (awareness/exposure, followed by extent of exposure and perpetrator identity if relevant), followed by questions about fears (including fear origins and subsequent behavioural changes if relevant). Next, participants answered questions about comfort with various forms of online participation, the psychological impact of online experiences, and engagement with safety tools online. At the end of the questions, participants were given an opportunity to provide feedback in a free text box before continuing to the debrief and finally completing the submission and being returned to Prolific for payment. The survey was designed to take approximately 15 minutes to complete and each participant received £2.25 for their time.
@@ 3.5 Data analysis
Additionally, we test for associations between fears about online harms and comfort engaging in several forms of online participation by running a series logistic regression analyses, each including fear as predictor (one for fear about each of 12 harms, see Results for harms included), and comfort engaging in the online behaviour of interest as the outcome (one for comfort with each of the seven behaviours), all controlling for gender, age, and internet use. For all statistical analyses, the accepted significance threshold was set at .05.
@@ 4.1.1 Awareness and exposure
Gender differences in exposure to various harms show mixed results. Women are 22% more likely to say they have seen online misogyny (p=.027), 24% more likely to say they have seen bullying (p=.017) and 104% more likely to say they have seen eating disorder promotion content (p<<.001) than men. However, women are 20% less likely to say they have seen hate speech (p=.016), 25% less likely to say they have seen misinformation (p=.017), 23% less likely to say they have seen seen impersonation (p=.007), 31% less likely to say they have seen threats of physical violence (p<<.001), and 29% less likely to say they have seen doxing (p=.003) than men. For many of the harms, such as trolling and cyberstalking, differences in self-reported exposure between men and women are not statistically significant.
@@
Image-based
sexual abuse
Female 8.4%
@@
Eating disorder
content
Female 27.1% ***
@@
Table 1: Proportions of men and women indicating they have seen each harm in the past year. In the Sig. column, * means significant at the p<<.05 level; ** at the p<<.01 level, and *** at the p<<.001 level.
@@ 4.1.2 Extent of exposure
Importantly, the results on exposure presented above show whether or not people have seen certain types of content but not whether they have been direct targets of online harms, nor how much. To understand the extent of people’s exposure more deeply, for each of the 15 harms that participants indicated they had seen in the past year, we analysed responses to follow up questions about extent to which they had witnessed such content and the extent to which they had directly received such content (see Methods). Response options for both scales were converted into 4 point numeric scales (1 = Never; 2 = Once; 3 = 2-4 times; 4 = 5 or more times, with ‘Prefer not to say’ as NA) and non-parametric t-tests (Mann-Whitney U-Tests) examined gender differences in being directly targeted by each type of harm. [^5] An alternative analysis is to create binary responses (Never = 0 / At least once = 1) and use logistic regressions to examine the effect of gender on whether people have been targeted before or not. Taking this approach yields identical results. If participants had previously indicated that they had not seen the harm in the past year and were therefore not asked follow-up questions about the capacity in which they had seen it, their responses for this question were re-coded from missing to ‘Never’.
@@
Women reported being the direct targets of several online harms to a significantly greater extent than men, including online misogyny (p<<.001), cyberstalking (p=.026), cyberflashing (p<<.001), eating disorder promotion content (p=.039) and, marginally, image based abuse (p=.075). Table 4 in Supplementary Information shows proportions of men and women choosing each response option for the five harms that women report being targeted by to a greater extent than men. Men reported being the direct targets of hate speech (p<<.001), misinformation (p<<.001), trolling (p<<.001) and threats of physical violence (p=.007) to a greater extent than women. Table 5 in Supplementary Information shows proportions of men and women choosing each response option for the four harms that men report being targeted by to a greater extent than women. There were no significant gender differences in reports of being directly targeted by the remaining six harms. Figure 1 shows the extent to which men and women report having been direct targets of each online harm. Note that while we asked individuals who had been targeted by harms whether they knew the identity(ies) of the perpetrator(s), comparing proportions of responses led to sample sizes too small to make meaningful comparisons for this question.
@@
Figure 1: The extent to which people report having been direct targets of each online harm by gender. F>>M indicates women are targeted significantly more than men. M<<F indicates men are targeted significantly more than women. ‘Prefer not to say’ responses were uncommon and are not included.
@@
Comparing gender differences in witnessing each type of harm (an individual seeing harmful content online not directly targeted at them), men reported witnessing hate speech, misinformation, trolling, impersonation, catfishing, threats of physical violence, threats of sexual violence, doxing, and image-based abuse to a greater extent than women (all ps<<.05), while women reported witnessing eating disorder promotion content (p<<.001) and, marginally, misogyny (p=.051) and bullying (p=.073) to a greater extent than men. [^6] We explored whether possible gender differences in overall internet use explained these results but we found men and women report using the internet and social media to a similar extent overall. The differences here in results for witnessing and directly receiving online harms shows the importance of understanding both who is most likely to be targeted, alongside understanding overall exposure.
@@ 4.2 Psychological impact of online experiences
We asked participants the extent to which experiences online had ever caused them six particular kinds of feelings (see Methods). For each of these responses, we created a binary outcome variable to indicate whether they had experienced this psychological effect or not (Yes, 1 - Somewhat/Very much and No, 0 - Not at all/Not very much). Overall, 48% of respondents said they had been left feeling sad or low as a result of an online experience; 68% that they had felt angry or frustrated; 11% with physical symptoms (such as insomnia, headaches and stomach aches); 9% as though their career had been negatively affected; 15% as though their relationships had been negatively affected; and 22% feeling (more) eager or keen to use social media to advocate for a specific cause. Logistic regression analyses with gender as predictor and response as outcome for each of the items found that women were significantly more likely than men to have been left feeling sad or low (96% more likely), angry or frustrated (54% more likely), and with physical symptoms (47% more likely) as a result of an experience online (all ps<<.001). There were no gender differences in responses for the remaining three questions about psychological impact (all ps >>.05). Table 6 in Supplementary Information shows responses to each item for men and women.
@@ 4.3.1 Overall extent of fears about receiving online harms
To compare gender differences in fears about exposure to each of the 15 online harms, we created a binary outcome for fear, with ‘Somewhat’ and ‘Very much’ as 1 - Fearful, and ‘Not at all’ and ‘Not very much’ as 0 - Not fearful. Women consistently express significantly greater levels of fear than men across all 15 types of content, both for witnessing and for directly receiving each harm. Women are 117% more likely to fear receiving hate speech than men, 39% more likely to fear receiving misinformation, 489% more likely to fear receiving misogyny, 125% more likely to fear being targeted by trolling and bullying, 120% more likely to fear being targeted by cyberstalking, 196% more likely to fear being targeted by cyberflashing, 68% more likely to fear being targeted by group attacks, 50% more likely to fear having an account impersonated, 64% more likely to fear catfishing, 112% more likely to fear receiving physically violent threats, 215% more likely to fear receiving sexually violent threats, 38% more likely to fear being targets of doxing, 133% more likely to fear being targets of image-based abuse, and 174% more likely to fear receiving eating disorder promotion content (all ps<<.001). Table 2 shows proportions of men and women fearing receiving each type of online harm.
@@
Image based
sexual abuse
Low fear 71.0% 51.2%
@@
Eating disorder
content
Low fear 87.3% 71.4%
@@ 4.3.2 Origins of fears about receiving online harms
When asked where they thought fears about receiving online harms might have come from, the top three choices for women were: The media (65%), Female friend’s experience (37%) and Public figure (35%). Additionally, 30% of women chose Personal experience, 8% chose Male friend’s experience, 5% Non-binary friend’s experience. The top three choices for men were: The media (60%), Personal experience (29%), and Female friend’s experience (27%). Additionally, 26% of men chose Public figure, 14% chose Male friend’s experience, and 3% Non-binary friend’s experience. Therefore, both men and women chose the media and female friends’ experiences as the most common sources of fears of online harms.
4.3.3 Behavioural effects of fears